# Learning to Drive Ethically: Moral Reasoning in Autonomous Driving

### v2

@@ Abstract
Autonomous vehicles hold great promise for reducing traffic fatalities and improving transportation efficiency, yet their widespread adoption hinges on embedding robust ethical reasoning into routine and emergency maneuvers, particularly to protect vulnerable road users (VRUs) such as pedestrians and cyclists. Here, we present a hierarchical Safe Reinforcement Learning (Safe RL) framework that explicitly integrates moral considerations with standard driving objectives. At the decision level, a Safe RL agent is trained using a composite ethical risk cost, combining collision probability and harm severity, to generate high-level motion targets. A dynamic Prioritized Experience Replay mechanism amplifies learning from rare but critical, high-risk events. At the execution level, polynomial path planning coupled with Proportional–Integral–Derivative (PID) and Stanley controllers translates these targets into smooth, feasible trajectories, ensuring both accuracy and comfort. We train and validate our approach on rich, real-world traffic datasets encompassing diverse vehicles, cyclists, and pedestrians, and demonstrate that it outperforms baseline methods in reducing ethical risk and maintaining driving performance. To our knowledge, this is the first study of ethical decision-making for autonomous vehicles via Safe RL evaluated on real-world, human-mixed traffic scenarios. Our results highlight the potential of combining formal control theory and data-driven learning to advance ethically accountable autonomy that explicitly protects those most at risk in urban traffic environments.
@@ 1 Introduction
Over the past several decades, both academic institutions and industry stakeholders have devoted significant effort to advancing autonomous driving (AD) technologies. By eliminating human error, autonomous vehicles (AVs) promise to reduce traffic accidents and enhance both social welfare and economic productivity. While fully functional Level 5 systems, with completely autonomous vehicles available on the mass market, are expected to emerge in the coming decades (Khan et al., 2022), the transitional period will involve extensive interaction between AVs and human-driven vehicles. This raises a pressing societal question: Can AVs make responsible decisions in emergencies and everyday interactions, not only to protect their occupants but also to protect those outside the vehicle (Othman, 2021)? This question is especially urgent for vulnerable road users (VRUs) such as pedestrians and cyclists, who bear a disproportionately high burden of severe injury and death in collisions. According to the U.S. National Highway Traffic Safety Administration, 7,522 pedestrians were killed and more than 67,000 injured in 2022 (Naumann, 2025). Safeguarding VRUs should therefore be a primary objective of technical design, regulation, and public policy (Campolettano et al., 2024). Beyond emergency responses, broader ethical and policy concerns have also emerged, including legal liability, data privacy, and equitable access to mobility solutions (Wang et al., 2020a; Lin, 2016; Bonnefon et al., 2016). Early public and academic debates often focused on abstract dilemmas such as the “trolley problem”, how an AV should react when a collision is unavoidable. Empirical work such as the Moral Machine Experiment has demonstrated that ethical preferences vary widely across cultures, for example, favoring younger over older pedestrians or humans over animals (Awad et al., 2018). However, emphasizing these rare, extreme scenarios risks obscuring the everyday ethical decisions AVs must make, such as maintaining safe following distances, navigating busy intersections, and yielding to pedestrians, which cumulatively determine public safety and perceived fairness in urban environments (Evans et al., 2020).
Recent research on embedding ethical reasoning into autonomous vehicles has increasingly combined formal control techniques with data-driven learning to develop hybrid ethical decision-making systems. For example, in (Wang et al., 2020b), a lexicographic risk-minimization framework is presented that ranks road users by priority and optimizes vehicle trajectories to reduce collision risk in descending order of that ranking. Similarly, (Geisslinger et al., 2023) proposes an ethical trajectory planning approach in which candidate paths are evaluated using a set of ethics metrics, total risk minimization, worst-off protection, equal treatment, liability-based responsibility, and bounded personal risk, and weighted according to regional guidelines. These control-theoretic methods leverage well-established optimization tools and offer verifiable safety guarantees. However, they often require careful parameter tuning and rely on pre-defined policies. Meanwhile, RL has gained traction as a promising approach for autonomous driving (Kiran et al., 2021; Li and Okhrin, 2023), thanks to its success in solving complex problems in domains such as game strategy and robotics control (Wurman et al., 2022; Mnih et al., 2015; Won et al., 2020; Li et al., 2024). Although many RL studies incorporate safety as a primary design principle by embedding safety constraints into reward functions, explicit ethical reasoning remains largely absent. To address this gap, autonomous vehicles should be capable of more than just reactive ethical choices in rare, high-risk situations. They should also anticipate ethical concerns in everyday scenarios to prevent crises before they emerge. This requires three key capabilities: (i) the ability to anticipate the intentions of surrounding traffic participants; (ii) the capacity to make moral decisions involving multiple agents; and (iii) consideration of broader driving objectives such as safety and efficiency.
To that end, we propose EthicAR, an ethical AD agent built upon the principles of Safe RL. EthicAR separates the objectives of AV into two components: standard driving goals (e.g., safety and efficiency) and moral reasoning across surrounding traffic participants. This structure allows the vehicle to make ethically informed decisions, particularly in the face of potential collisions. The framework is organized into two hierarchical levels: a decision level and an execution level. The decision level does not directly control the vehicle, instead, it determines how and where the vehicle should move, producing a target lateral position and target longitudinal speed. It is implemented using a Safe RL algorithm that incorporates ethical risk as a cost signal, following the formulation introduced in (Geisslinger et al., 2023). The risk metric combines two factors: (i) the probability of collision involving either the ego vehicle or other road users, and (ii) the severity of harm associated with those collisions. Because high-risk situations are rare in naturalistic driving data, we introduce a dynamic Prioritized Experience Replay (PER) mechanism. This strategy emphasizes learning from high-risk experiences, thereby enhancing the ability to learn from rare but critical situations. The execution level translates high-level decisions into physical motion. It includes a path planner that generates smooth, feasible trajectories using polynomial equations. A low-level path follower then executes these trajectories using a PID controller for longitudinal speed and a Stanley controller for lateral position tracking (Thrun et al., 2006). This setup ensures stable and comfortable vehicle behavior. To train the agent, we use real-world datasets that capture diverse traffic scenarios with various types of vehicles, pedestrians, and cyclists. These datasets provide realistic interactions, allowing the learned policy to generalize effectively to heterogeneous and dynamic environments. We evaluate EthicAR against several baselines in real-world scenarios and present a comprehensive analysis of its performance. To the best of our knowledge, this is the first study to apply Safe RL to ethical decision-making in AVs using real-world data, backed by extensive empirical validation. In summary, the main contributions of this paper are as follows:
@@
    Ethics-Aware Cost Function for Safe RL. We leverage a Safe RL framework to balance standard driving objectives with ethical decision-making. By integrating an ethics-inspired risk metric into the cost function, the agent is encouraged not only to satisfy normal driving requirements but also to minimize the overall risk posed to all road users.
@@
    Dynamic PER for Rare Events. To improve learning from rare but high-risk situations, we introduce a dynamic PER mechanism. This approach increases the sampling frequency of critical experiences during training. We also conducted an ablation study to demonstrate its effectiveness.
@@
    Two-Level Control Architecture. Our framework adopts a hierarchical structure consisting of a decision level and an execution level. The decision level generates high-level motion targets using Safe RL, while the execution level ensures smooth and comfortable vehicle motion through a polynomial path planner and PID/Stanley controllers.
@@
    Comprehensive Evaluation in Real-World Scenarios. We evaluate the proposed framework against baselines using real-world traffic datasets. Our analysis includes quantitative and qualitative comparisons from multiple perspectives, demonstrating the practical value of our approach.
Figure 1: Overall diagram of the proposed data-driven EthicAR methodology. The hierarchical architecture comprises a decision level, where a Safe RL agent with an ethics-aware cost function combining collision probability and harm severity, learns to output target motions. A dynamic PER emphasizes rare, high-risk events during training. The execution level then converts these high-level commands into smooth, feasible trajectories via a polynomial path planner and tracks them using PID/Stanley controllers, ensuring safe, comfortable maneuvers.
@@ 2 Background
RL is a foundational component of artificial intelligence, where an agent learns to make decisions through trial-and-error interactions with its environment (Sutton and Barto, 2018). This learning process is typically formalized as a Markov Decision Process (MDP), represented by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$. Here, $\mathcal{S}$ denotes the state space, capturing all possible configurations of the environment. $\mathcal{A}$ denotes the action space, describing the set of actions available to the agent. The transition probability function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ models the likelihood of transitioning from one state to another, given an action. The reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns scalar feedback to each state-action pair. The discount factor $\gamma\in[0,1)$ determines the relative importance of future rewards compared to immediate ones. Within this framework, the goal of the agent is to learn a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, which maps states to a probability distribution over actions, so as to maximize the expected cumulative discounted return:
@@
To assess the effectiveness of a policy $\pi$, two value functions are commonly used. The state-value function, $V^{\pi}(s_{t})=\mathbb{E}_{\pi}\left[R_{t}\mid s_{t}\right]$, estimates the expected return when starting in state $s_{t}$ and following policy $\pi$ thereafter. The action-value function, $Q^{\pi}(s_{t},a_{t})=\mathbb{E}_{\pi}\left[R_{t}\mid s_{t},a_{t}\right]$, measures the expected return of taking action $a_{t}$ in state $s_{t}$, then following $\pi$. Based on how these value functions are utilized, RL methods are broadly categorized into two groups: value-based and policy gradient methods. In value-based approaches, an optimal policy $\pi^{\star}$ is derived by acting greedily to the optimal action-value function $Q^{\star}$. This optimal function satisfies the condition $Q^{\star}(s,a)=\max_{\pi}Q^{\pi}(s,a)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Conversely, policy gradient methods aim to directly optimize a parameterized policy $\pi_{\theta}$ by maximizing a performance objective, such as the expected return under $\pi_{\theta}$. A prominent subclass of these methods is the actor-critic architecture, where the actor represents the policy and the critic estimates value functions to guide the updates of the actor (Haarnoja et al., 2018).
@@
Safe RL extends the RL paradigm by explicitly incorporating safety constraints into the learning process. This is formalized using a Constrained Markov Decision Process (CMDP), which augments the classic MDP tuple with an additional cost function $\mathcal{C}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ and a predefined safety threshold $\eta$ Altman (2021). The cost function $\mathcal{C}(s,a)$ quantifies the risk or "unsafety" of taking action $a$ in state $s$. The expected cumulative cost under a given policy $\pi$ is defined as: $D({\pi})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{C}(s_{t},a_{t},s_{t+1})\right]$. A policy is considered feasible if it satisfies the safety constraint, i.e., $D(\pi)\leq\eta$. The set of all such feasible policies is denoted as $\Pi_{\mathcal{C}}=\left\{\pi\in\Pi|\mathcal{D}(\pi)\leq\eta\right\}$. The objective in Safe RL is to maximize the expected return while ensuring the cumulative cost remains within the allowed threshold. This leads to the following constrained optimization problem:
@@
where $\pi_{\theta}$ is a parametrized policy. To solve such constrained problems, Lagrangian relaxation (Boyd and Vandenberghe, 2004) and projection-based policy updates (Achiam et al., 2017) are commonly used. The Lagrangian method introduces a non-negative dual variable $\lambda\geq 0$ to convert the constrained problem into an unconstrained saddle-point formulation:
@@ 3.2 LSTM based SACLag
Soft Actor-Critic (SAC) algorithm combines actor-critic architecture with entropy regularization. It aims to maximize both cumulative rewards and policy entropy, encouraging exploration by favoring stochastic policies. SAC employs a temperature parameter $\alpha$ to balance reward maximization and entropy trade-offs. This temperature is automatically adjusted to maintain a target entropy level, avoiding manual tuning. The actor optimizes a stochastic policy, while the critic estimates Q-values using clipped double-Q networks to reduce overestimation bias (Haarnoja et al., 2018). SAC with Lagrangian relaxation (SACLag) uses a Lagrangian multiplier to optimize while enforcing the cost constraint dynamically. It transforms the problem into a dual optimization: the primal variables (policy and Q-networks) maximize the reward-entropy objective, while the dual variable (Lagrangian multiplier) is adjusted via gradient ascent on the Lagrange loss.
As outlined in Section 3.1, the objective of the RL agent is to dynamically supply the path planner with relevant information for trajectory generation. However, individual states frequently lack the contextual depth needed to support optimal decision-making by the agent. In particular, historical information about the ego vehicle and surrounding traffic is important. To formally account for this, we model the problem as a partially observable Markov decision process (POMDP) (Kaelbling et al., 1998), which extends the MDP tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$ with an observation space $\mathcal{O}$ and an observation function $\mathcal{Z}:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{O})$, where $\Delta(\mathcal{O})$ denotes the probability distribution over $\mathcal{O}$. In a POMDP, the agent does not directly observe the true state $s_{t}\in\mathcal{S}$ but instead receives an observation $o_{t}\in\mathcal{O}$ generated according to $\mathcal{Z}(o_{t}\mid s_{t},a_{t-1})$. To address this partial observability, we integrate Long Short-Term Memory (LSTM) networks (Hochreiter and Schmidhuber, 1997) into the SACLag framework to deal with POMDPs. Each network in this framework incorporates LSTM layers, which process sequences of past time steps. This temporal reasoning capability enables the agent to retain and exploit historical information in combination with the current observation, thereby supporting more informed and consistent decision-making.

### v3

# Ethics-Aware Safe Reinforcement Learning for Rare-Event Risk Control in Interactive Urban Driving

@@ Abstract
Autonomous vehicles hold great promise for reducing traffic fatalities and improving transportation efficiency, yet their widespread adoption hinges on embedding credible and transparent ethical reasoning into routine and emergency maneuvers, particularly to protect vulnerable road users (VRUs) such as pedestrians and cyclists. Here, we present a hierarchical Safe Reinforcement Learning (Safe RL) framework that augments standard driving objectives with ethics-aware cost signals. At the decision level, a Safe RL agent is trained using a composite ethical risk cost, combining collision probability and harm severity, to generate high-level motion targets. A dynamic, risk-sensitive Prioritized Experience Replay mechanism amplifies learning from rare but critical, high-risk events. At the execution level, polynomial path planning coupled with Proportional-Integral–Derivative (PID) and Stanley controllers translates these targets into smooth, feasible trajectories, ensuring both accuracy and comfort. We train and validate our approach on closed-loop simulation environments derived from large-scale, real-world traffic datasets encompassing diverse vehicles, cyclists, and pedestrians, and demonstrate that it outperforms baseline methods in reducing risk to others while maintaining ego performance and comfort. This work provides a reproducible benchmark for Safe RL with explicitly ethics-aware objectives in human-mixed traffic scenarios. Our results highlight the potential of combining formal control theory and data-driven learning to advance ethically accountable autonomy that explicitly protects those most at risk in urban traffic environments. Across two interactive benchmarks and five random seeds, our policy decreases conflict frequency by $25\sim 45\%$ compared to matched task successes while maintaining comfort metrics within 5%.
@@ 1 Introduction
Over the past several decades, both academic institutions and industry stakeholders have devoted significant effort to advancing autonomous driving (AD) technologies. By eliminating human error, autonomous vehicles (AVs) promise to reduce traffic accidents and enhance both social welfare and economic productivity. While fully functional Level 5 systems are still some decades away [khan2022level], the transitional period will involve extensive interaction between AVs and human-driven vehicles. This raises a pressing societal question: Can AVs make responsible decisions in emergencies and everyday interactions, not only to protect their occupants but also to protect those outside the vehicle [othman2021public]? This question is especially urgent for vulnerable road users (VRUs) such as pedestrians and cyclists, who bear a disproportionately high burden of severe injury and death in collisions. According to the U.S. National Highway Traffic Safety Administration, 7,522 pedestrians were killed and more than 67,000 injured in 2022 [naumann2025pedestrian]. Safeguarding VRUs should therefore be a primary objective of technical design, regulation, and public policy [campolettano2024baseline]. Beyond emergency responses, broader ethical and policy concerns have also emerged, including legal liability, data privacy, and equitable access to mobility solutions [wang2020ethical, lin2016ethics, bonnefon2016social]. Early debates often focused on abstract dilemmas such as the “trolley problem”. Empirical work such as the Moral Machine Experiment has demonstrated that ethical preferences vary widely across cultures [awad2018moral]. However, focusing on such extreme dilemmas risks obscuring the everyday ethical trade-offs AVs face, such as yielding to pedestrians at crosswalks, balancing risk across multiple agents, and ensuring equitable safety for VRUs. These frequent, small-scale interactions cumulatively determine both public safety and social trust in autonomy [evans2020ethical].
Recent research has begun embedding ethical reasoning into AVs by combining formal control techniques with data-driven learning. For example, [wang2020ethical2] introduces a lexicographic risk-minimization framework that ranks road users by priority, while [geisslinger2023ethical] proposes an ethical trajectory planning approach in which candidate paths are evaluated using a set of principles such as risk minimization, worst-off protection, and equal treatment. These approaches leverage optimization and verification tools, offering interpretability and theoretical guarantees. At the same time, they often rely on pre-defined rankings, hand-tuned parameters, or fixed weighting schemes, which may not adapt well to heterogeneous, real-world traffic. In parallel, Reinforcement Learning (RL) has gained traction in AD research [kiran2021deep, li2023modified], drawing on its success in high-dimensional control problems [mnih2015human, wurman2022outracing, li2025autonomous]. RL methods can adapt through data and explore trade-offs automatically. Still, most RL-based driving studies have focused on safety constraints or reward shaping, without explicitly operationalizing ethical reasoning across multiple road users. This gap limits their ability to address fairness and accountability in shared traffic environments.
To bridge this gap, we propose EthicAR, an ethics-aware Safe RL framework that integrates moral considerations as a continuous design principle rather than an afterthought for rare emergencies. EthicAR separates the objectives of AVs into two components: standard driving goals (e.g., safety and efficiency) and ethics-aware costs across surrounding traffic participants. The framework is organized into two hierarchical levels: a decision level, where a Safe RL algorithm uses ethics-inspired costs to select high-level motion targets; and an execution level, where classical controllers ensure smooth, feasible, and comfortable trajectory following. To better learn from rare but critical events, we introduce a dynamic, risk-sensitive Prioritized Experience Replay mechanism. We train and evaluate EthicAR in closed-loop simulation environments derived from large-scale, real-world traffic datasets with heterogeneous road users, enabling reproducible and realistic evaluation across complex scenarios.
The contributions of this paper are as follows:
@@
    Ethics-Aware Safe RL. We design a composite ethical risk cost that integrates collision probability and harm severity, enabling Safe RL to balance ego safety with fairness to other road users.
@@
    Dynamic PER for Rare Events. We introduce a risk-sensitive replay scheme that improves learning efficiency in rare but safety-critical interactions, validated through ablation.
@@
    Two-Level Control Architecture. Our framework separates high-level decision making (Safe RL) from low-level execution (classical controllers), ensuring both principled reasoning and stable motion.
@@
    Comprehensive Evaluation. We benchmark EthicAR against state-of-the-art baselines in simulation environments derived from real-world data, with stratified analysis across diverse traffic scenarios.
Limitations. While this study shows the potential of ethics-aware Safe RL for AVs, several limitations remain. First, our evaluation relies on dataset-derived closed-loop simulations, where surrounding agents are only partially reactive; fully interactive multi-agent dynamics require further validation. Second, the ethical cost function depends on chosen weights and harm models, which may reflect implicit normative assumptions; systematic calibration or participatory preference elicitation remains future work. Third, our probabilistic risk model assumes independence across timesteps, which may underestimate compounding risk. Finally, we do not claim formal safety guarantees; rather, we aim to provide a reproducible benchmark and a step toward more accountable autonomous decision-making.
Figure 1: Overall diagram of the proposed data-driven EthicAR methodology. The hierarchical architecture comprises a decision level, where a Safe RL agent with an ethics‐aware cost function combining collision probability and harm severity, learns to output target motions. A dynamic PER emphasizes rare, high‐risk events during training. The execution level then converts these high‐level commands into smooth, feasible trajectories via a polynomial path planner and tracks them using PID/Stanley controllers, ensuring safe, comfortable maneuvers.
@@ 2 Background
RL is a foundational component of artificial intelligence, where an agent learns to make decisions through trial-and-error interactions with its environment [Sutton2018]. This learning process is typically formalized as a Markov Decision Process (MDP), represented by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$. Here, $\mathcal{S}$ denotes the state space, capturing all possible configurations of the environment. $\mathcal{A}$ denotes the action space, describing the set of actions available to the agent. The transition probability function $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ models the likelihood of transitioning from one state to another, given an action. The reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns scalar feedback to each state-action pair. The discount factor $\gamma\in[0,1)$ determines the relative importance of future rewards compared to immediate ones. Within this framework, the goal of the agent is to learn a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, which maps states to a probability distribution over actions, so as to maximize the expected cumulative discounted return:
@@
To assess the effectiveness of a policy $\pi$, two value functions are commonly used. The state-value function, $V^{\pi}(s_{t})=\mathbb{E}_{\pi}\left[R_{t}\mid s_{t}\right]$, estimates the expected return when starting in state $s_{t}$ and following policy $\pi$ thereafter. The action-value function, $Q^{\pi}(s_{t},a_{t})=\mathbb{E}_{\pi}\left[R_{t}\mid s_{t},a_{t}\right]$, measures the expected return of taking action $a_{t}$ in state $s_{t}$, then following $\pi$. Based on how these value functions are utilized, RL methods are broadly categorized into two groups: value-based and policy gradient methods. In value-based approaches, an optimal policy $\pi^{\star}$ is derived by acting greedily to the optimal action-value function $Q^{\star}$. This optimal function satisfies the condition $Q^{\star}(s,a)=\max_{\pi}Q^{\pi}(s,a)$ for all $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Conversely, policy gradient methods aim to directly optimize a parameterized policy $\pi_{\theta}$ by maximizing a performance objective, such as the expected return under $\pi_{\theta}$. A prominent subclass of these methods is the actor-critic architecture, where the actor represents the policy and the critic estimates value functions to guide the updates of the actor [haarnoja2018soft].
@@
Safe RL extends the RL paradigm by explicitly incorporating safety constraints into the learning process. This is formalized using a Constrained Markov Decision Process (CMDP), which augments the classic MDP tuple with an additional cost function $\mathcal{C}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ and a predefined safety threshold $\eta$ [altman2021constrained]. The cost function $\mathcal{C}(s,a)$ quantifies the risk or "unsafety" of taking action $a$ in state $s$. The expected cumulative cost under a given policy $\pi$ is defined as: $D({\pi})=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{C}(s_{t},a_{t},s_{t+1})\right]$. A policy is considered feasible if it satisfies the safety constraint, i.e., $D(\pi)\leq\eta$. The set of all such feasible policies is denoted as $\Pi_{\mathcal{C}}=\left\{\pi\in\Pi|\mathcal{D}(\pi)\leq\eta\right\}$. The objective in Safe RL is to maximize the expected return while ensuring the cumulative cost remains within the allowed threshold. This leads to the following constrained optimization problem:
@@
where $\pi_{\theta}$ is a parametrized policy. To solve such constrained problems, Lagrangian relaxation [boyd2004convex] and projection-based policy updates [achiam2017constrained] are commonly used. The Lagrangian method introduces a non-negative dual variable $\lambda\geq 0$ to convert the constrained problem into an unconstrained saddle-point formulation:
@@ 3.2 LSTM based SACLag
Soft Actor-Critic (SAC) algorithm combines actor-critic architecture with entropy regularization. It aims to maximize both cumulative rewards and policy entropy, encouraging exploration by favoring stochastic policies. SAC employs a temperature parameter $\alpha$ to balance reward maximization and entropy trade-offs. This temperature is automatically adjusted to maintain a target entropy level, avoiding manual tuning. The actor optimizes a stochastic policy, while the critic estimates Q-values using clipped double-Q networks to reduce overestimation bias [haarnoja2018soft]. SAC with Lagrangian relaxation (SACLag) uses a Lagrangian multiplier to optimize while enforcing the cost constraint dynamically. It transforms the problem into a dual optimization: the primal variables (policy and Q-networks) maximize the reward-entropy objective, while the dual variable (Lagrangian multiplier) is adjusted via gradient ascent on the Lagrange loss.
As outlined in Section 3.1, the objective of the RL agent is to dynamically supply the path planner with relevant information for trajectory generation. However, individual states frequently lack the contextual depth needed to support optimal decision-making by the agent. In particular, historical information about the ego vehicle and surrounding traffic is important. To formally account for this, we model the problem as a partially observable Markov decision process (POMDP) [kaelbling1998planning], which extends the MDP tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$ with an observation space $\mathcal{O}$ and an observation function $\mathcal{Z}:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{O})$, where $\Delta(\mathcal{O})$ denotes the probability distribution over $\mathcal{O}$. In a POMDP, the agent does not directly observe the true state $s_{t}\in\mathcal{S}$ but instead receives an observation $o_{t}\in\mathcal{O}$ generated according to $\mathcal{Z}(o_{t}\mid s_{t},a_{t-1})$. To address this partial observability, we integrate Long Short-Term Memory (LSTM) networks [hochreiter1997long] into the SACLag framework to deal with POMDPs. Each network in this framework incorporates LSTM layers, which process sequences of past time steps. This temporal reasoning capability enables the agent to retain and exploit historical information in combination with the current observation, thereby supporting more informed and consistent decision-making.